evening shift summary 6/7/99

AG (ag@fnal.gov)
Tue, 8 Jun 1999 01:00:06 -0400 (EDT)

On Shift: Rick, Sasha

Hours of beam delivered at all intensities: ~ 3.
Hours of beam delivered at nominal intensity: ~ 3.
Hours of E832 data (special runs not included): 0.
Hours of special runs (laser scans, muon runs...): ~ 1.

Accounting of lost time: waiting for the NM3S to be fixed.

========================================================================

Runs ended this shift:

13421 - Long Laser Scan
13422 - Cosmics
13425 - BEAM_832_ECAL ( in progress )

Tapes brought to FCC this shift: 0

========================================================================

Accesses during shift:

Hogan and Sasha made access. Hogan increased laser light level by slightly
less than 2 times.

========================================================================
Description of Shift:

The goal of the shift was to collect 2 million of ECAL triggers for CsI,
at the end of this shift we have 1.4 million and the run is still in
progress.

DETAILS of SHIFT:
------------------

Rick re-made L3 drivers with a few new histos to monitor beam structure.

BEAM_MU_4E11 trigger modified to include B08=Accid 90 so that we can
monitor beam structure with mu-runs, using new online plots 3p/trg/401 &
3p/trg/402.

Ed updated reference for pmt gain drift plot. Old version renamed to
eopd_refitxt_Jun7.

At 21:05 beam stop crush button in ktev control room was tested.
Hitting switch closed beam stop, but did not drop primary beam permit.
Opening switch left beam stop closed.

Peter performed calibration and installed new dpmt constants (12425 -
99999).

At 21:10 we've got beam. Start run 13425 ( BEAM_832_ECAL),
initially requesting 4E11 in 6 bunches, to perform calibrations.

Rick notices that MA pedestal smears on hadrons, although it looked ok on
muons.